### v2

# Reasoning Models Better Express Their Confidence

Abstract
Despite their strengths, large language models (LLMs) often fail to communicate their confidence accurately, making it difficult to assess when they might be wrong and limiting their reliability. In this work, we demonstrate that reasoning models that engage in extended chain-of-thought (CoT) reasoning exhibit superior performance not only in problem-solving but also in accurately expressing their confidence. Specifically, we benchmark six reasoning models across six datasets and find that they achieve strictly better confidence calibration than their non-reasoning counterparts in 33 out of the 36 settings. Our detailed analysis reveals that these gains in calibration stem from the slow thinking behaviors of reasoning models (e.g., exploring alternative approaches and backtracking) which enable them to adjust their confidence dynamically throughout their CoT, making it progressively more accurate. In particular, we find that reasoning models become increasingly better calibrated as their CoT unfolds, a trend not observed in non-reasoning models. Moreover, removing slow thinking behaviors from the CoT leads to a significant drop in calibration. Lastly, we show that non-reasoning models also demonstrate enhanced calibration when simply guided to slow think via in-context learning, fully isolating slow thinking as the source of the calibration gains. [^1] Our code is available at https://github.com/MattYoon/reasoning-models-confidence
Question Who’s the Athenian reformer that set up the council of 500?
Model CoT … Hmm, I remember that Athens had several important reformers, like Solon, Cleisthenes, and Pericles. … So maybe it wasn’t Solon. Then there’s Cleisthenes. … (1) Oh, and he also created the council of 500, which was … (2) Pericles came later, and he was … (3) but I don’t think he was the one who established the council of 500. … (4) So putting it all together, Cleisthenes is the reformer who … I don’t think I’m mixing up any details here.
[…]
1 Introduction
A persistent weakness of large language models (LLMs) is their tendency to sound confident even when they are wrong (huang2025survey, ; xiong2024can, ). This overconfidence threatens their reliability, especially in high-stakes scenarios (pal2023med, ; zhang2023siren, ). Meanwhile, recent reasoning models like OpenAI’s o1 (jaech2024openai, ) and Deepseek-R1 (deepseekr1, ) have demonstrated strong problem-solving capabilities through chain-of-thought (CoT) (wei2022chain, ) eliciting slow thinking behaviors such as exploring alternative approaches and verifying their answers (min2024imitate, ; gandhi2025cognitive, ). Yet, it remains underexplored whether such slow thinking behaviors also help LLMs to “know what they know” (kadavath2022language, ) and accurately communicate the limits of their knowledge.
To this end, we present an extensive study of reasoning models’ ability to accurately estimate and express confidence in their output; a process known as verbalized confidence estimation (tian-etal-2023-just, ; geng-etal-2024-survey, ). We demonstrate that reasoning models exhibit superior confidence calibration compared to non-reasoning models, and that this improvement stems from slow thinking behaviors, which allow LLMs to dynamically adjust their confidence throughout the reasoning process. For instance, as shown in Figure 1, the model’s confidence increases when it first verifies its answer (“Oh, and he also created the council of 500”), and decreases when it considers an alternative (“Pericles came later”). The confidence then rises again as it rejects the alternative (“but I don’t think he was”), and reaches its highest level with further verification (“I don’t think I’m mixing up any details”).
We begin by benchmarking six reasoning models derived from four backbone LLMs against their non-reasoning, instruction-tuned counterparts across two knowledge-focused datasets (TriviaQA and NonambigQA) (joshi-etal-2017-triviaqa, ; min-etal-2020-ambigqa, ; kwiatkowski-etal-2019-natural, ) and four reasoning-intensive datasets (subsets of SuperGPQA and MMLU-Pro) (mmlupro, ; supergpqa, ). In 33 out of 36 settings, reasoning models strictly outperform their non-reasoning counterparts across all measured calibration metrics. Notably, reasoning models achieve better calibration even on knowledge-focused datasets, despite having comparable task accuracy to non-reasoning models. This highlights that the calibration gains are not simply derived from superior task performance.
We then conduct a detailed analysis to support our claim that the calibration gains in reasoning models stem from slow thinking, which allows them to dynamically adjust and correct their confidence throughout the reasoning process. (1) We measure changes in calibration metrics throughout the CoT process and observe a steady, gradual improvement for reasoning models, with statistically significant trends ($p<0.05$). Non-reasoning models show no significant trend and in some cases, surprisingly, even exhibit a worsening pattern. (2) In our ablation study, we find that the non-linear structure of slow thinking gandhi2025cognitive , such as the model’s ability to explore alternatives and revise its reasoning, plays a critical role in improving calibration. Conversely, prompting the model to explicitly reason about its own confidence yields only limited gains in calibration. (3) Lastly, we observe that non-reasoning models can also benefit from slow thinking when simply guided via in-context learning (dong2022survey, ), supporting our claim that these improvements arise from the slow thinking process itself, rather than from some other factors inherent to reasoning models.
2 Related Work
Confidence in LLMs
Confidence refers to the model’s estimated probability that its answer is correct, while calibration measures how well this confidence aligns with actual correctness (guo2017calibration, ). In other words, a well-calibrated model’s confidence is a reliable indicator of prediction correctness (nixon2019measuring, ). Therefore, assessing an LLM’s confidence and enhancing calibration enables more reliable and trustworthy models (geng-etal-2024-survey, ), supporting a wide range of applications such as hallucination detection (manakul-etal-2023-selfcheckgpt, ; varshney2023stitch, ), ambiguity detection (hou2024decomposing, ), uncertainty-guided data retrieval (jiang-etal-2023-active, ), and uncertainty-aware LLM agents (han-etal-2024-towards, ).
Various approaches have been proposed for confidence estimation, which is the process of inferring an LLM’s confidence in its predictions (geng-etal-2024-survey, ). This includes measuring token-level probabilities of answers (gupta2024language, ), training proxy probe models on internal hidden states (kadavath2022language, ), and applying supervised fine-tuning with labeled confidence data (xu2024sayself, ). Nonetheless, these methods assume access to a model’s internal states or weights, which are usually unavailable in proprietary LLMs (shrivastava2023llamas, ; xie2024survey, ). Additionally, they are often model-specific, necessitating readjustment or retraining when the underlying LLM changes (yang2024verbalized, ).
An alternative approach that avoids these limitations involves sampling multiple responses for the same prompt and deriving confidence based on response consistency (manakul-etal-2023-selfcheckgpt, ). While this method is compatible with proprietary LLMs and is model-agnostic, its primary limitation is the increased computational cost due to repeated inference (yang2024verbalized, ).
In this work, we focus on verbalized confidence estimation, where the LLM is prompted to directly express its confidence as a part of their output, either through a linguistic phrase (e.g., “Highly likely”) or a numerical probability (e.g., “0.95”) (geng-etal-2024-survey, ). Verbalized confidence estimation has emerged as a widely studied approach as it is compatible with proprietary LLMs, model-agnostic, and computationally efficient (tian-etal-2023-just, ; xiong2024can, ; yang2024verbalized, ). While verbalized confidence estimation offers ease of use and broad applicability, its accessibility also makes its limitations more consequential. Overconfidence is a persistent issue: LLMs tend to express high certainty even when they are incorrect, posing serious risks for real-world deployment of LLMs (xiong2024can, ; zhou2024relying, ). Our results suggest that reasoning models and their use of slow thinking hold strong potential to mitigate this issue.
Reasoning Models
Increasing compute at inference time, or inference-time scaling, has become a focus for improving the reasoning performance of LLMs (snell2025scaling, ). One effective approach to inference-time scaling is training models to generate longer CoTs, as demonstrated by models like OpenAI’s o1 (jaech2024openai, ) and DeepSeek-R1 (deepseekr1, ), which are commonly referred to as reasoning models. The key factor that distinguishes the CoT of reasoning models from that of non-reasoning, instruction-tuned models beyond just length is their slow thinking (min2024imitate, ), which drives their problem-solving ability. Slow thinking mirrors human cognitive behaviors, characterized by non-linear reasoning traces such as exploring alternative approaches, verifying answers, and backtracking (gandhi2025cognitive, ). During slow thinking, reasoning models also frequently produce epistemic markers that signal uncertainty (e.g., “I think”, “maybe”) which are rarely observed in non-reasoning models (zhou2024relying, ). Structurally, reasoning models typically enclose their CoT within special tokens such as <think>...</think> to isolate the reasoning phase, in contrast to non-reasoning models.
In this paper, we investigate whether reasoning models can more accurately express their confidence by leveraging their slow thinking behaviors. A concurrent study also examines the calibration of reasoning models (zhang2025reasoning, ), but focuses on training an external probe over hidden states to optimize CoT generation. In contrast, our study centers on verbalized confidence estimation and provides extensive analysis to uncover how reasoning models are able to better express their confidence.
3 Reasoning models better express their confidence
In this section, we demonstrate that reasoning models express their confidence more accurately than non-reasoning models, first detailing the experimental setup (Section 3.1) and then presenting the results (Section 3.2). Full results and additional experiments are provided in Appendix A, and further setup details are provided in Appendix B.
3.1 Experiment setup
Datasets
We evaluate LLM calibration across two types of datasets: knowledge-focused and reasoning-intensive. We use the knowledge-focused datasets, TriviaQA and NonambigQA (joshi-etal-2017-triviaqa, ; min-etal-2020-ambigqa, ; kwiatkowski-etal-2019-natural, ), to provide a more controlled setting for comparing the calibration of reasoning and non-reasoning models, as both achieve similar accuracy given that CoT does little to help solve these tasks (sprague2024cot, ). Therefore, this setup ensures that any differences in calibration are not simply due to reasoning models being better at solving the task. In addition, since reasoning models are primarily intended for complex reasoning tasks, we also include reasoning-intensive datasets, MMLU-Pro and SuperGPQA (mmlupro, ; supergpqa, ), to ensure our findings generalize. These datasets are challenging multiple-choice benchmarks centered on knowledge-driven reasoning. We use two subsets of each reasoning dataset: a Math subset focused on arithmetic reasoning, and a Non-Math subset covering other types of reasoning. Due to our broad range of experiments, we uniformly sample 1,000 examples from each dataset or subset to keep the compute manageable. To assess the variability of this choice, we perform bootstrapping over multiple resampled subsets and report the standard deviation in Appendix A.2.4.
Models
We benchmark a diverse set of six reasoning models derived from four different 32B-scale LLMs against their non-reasoning counterparts. [^2] We discuss the results on different of model sizes in Section 5. Specifically, we evaluate: R1-Distill-Qwen, QwQ, and OR1-Preview reasoning models (deepseekr1, ; qwq32b, ; skywork-or1-2025, ) against Qwen2.5-Instruct (qwen25technicalreport, ); GLM-Z1-0414 against GLM-4-0414 (glm2024chatglm, ); EXAONE-Deep (exaonedeep, ) against EXAONE-3.5-Instruct (exaone35, ); and Qwen3 with Thinking Mode (abbreviated as Qwen3 Thinking) against Qwen3 with Non-thinking Mode (Qwen3 Non-thinking) (qwen3, ). Unlike the other pairs, which consist of separate model checkpoints, Qwen3 is a hybrid model that supports both reasoning-style CoT (Thinking Mode) and non-reasoning-style CoT (Non-Thinking Mode) via prompting. [^3] The Non-thinking mode is enabled by injecting an empty thinking block, <think></think>, at the beginning of the model’s response.
Inference procedure
In a single turn of conversation, we instruct the models to perform three steps using CoT: (1) Solution Reasoning, where the model reasons step-by-step to arrive at an answer to the given question; (2) Confidence Reasoning, where it evaluates its own confidence for that answer step-by-step; and (3) Confidence Verbalization, where it maps their confidence in one of ten bins, ranging from “Almost no chance (0–0.1)” to “Almost certain (0.9–1.0)”. Each bin includes both a linguistic descriptor (e.g., “Almost certain”) and its corresponding numerical probability (e.g., “0.9–1.0”), which are inspired by the approaches used in prior work (tian-etal-2023-just, ; xiong2024can, ). To ensure a fair comparison, we use the same instructions for both reasoning and non-reasoning models. The full prompt we use is included in Appendix B.2.
For reasoning models, we expect all three steps to be carried out within the thinking process <think>...</think>. However, we observe that some reasoning models (specifically R1-Distill, OR1-Preview, and GLM-Z1) rarely engage in Confidence Reasoning. [^4] To elaborate, their thinking process mostly only contains Solution Reasoning, although they consistently do output their Confidence Verbalization outside of the thinking termination token </think> without CoT. To address this, we force these models to include Confidence Reasoning within their thinking process by generating up to the </think> token, and replacing the token with “Okay, now let’s assess my overall thinking process so far step-by-step. I need to evaluate how likely my answer is correct.” and performing a second round of inference. We ablate the effect of this choice in Section 4.2, where we observe that it has negligible impact on calibration.
As part of the instruction, models are asked to format their final response as “Answer:ANSWER Confidence:CONFIDENCE” after completing all three steps. We apply rule-based filtering using regular expressions to extract the predicted answer and confidence. Answers are matched against the ground truth using GPT-4o mini (openai_gpt4omini, ), leveraging the prompt from OpenAI’s simple-evals codebase (openai_simpleevals, ). We describe the full procedure we use to ensure that all model responses contain extractable answers and confidence predictions in Appendix B.2.
Table 1: Benchmark result on knowledge-focused datasets. Reasoning models are highlighted in blue, and the best performance within each backbone LLM group is shown in bold. Lower ECE and Brier Scores, and higher AUROC, indicate better calibration. Accuracy is included for reference.
TriviaQA NonambigQA
Model Acc. ECE $\downarrow$ Brier $\downarrow$ AUROC $\uparrow$ Acc. ECE $\downarrow$ Brier $\downarrow$ AUROC $\uparrow$
Qwen2.5-32B
Qwen2.5-Inst. 0.718 0.129 0.176 0.769 0.517 0.297 0.303 0.720
[…]
GLM-4-32B-0414
GLM-4-0414 0.814 0.084 0.137 0.675 0.640 0.246 0.269 0.643
[…]
EXAONE-3.5-32B
EXAONE-3.5-Inst. 0.715 0.130 0.178 0.749 0.511 0.302 0.302 0.721
[…]
Qwen3-32B
Qwen3 Non-thinking 0.711 0.207 0.230 0.650 0.511 0.403 0.403 0.572
[…]
Evaluation metrics
We use commonly adopted calibration metrics from prior work (geng-etal-2024-survey, ; tian-etal-2023-just, ; xiong2024can, ). Expected Calibration Error (ECE(naeini2015obtaining, ) measures the average discrepancy between accuracy and predicted confidence within each confidence bin, weighted by the number of samples in each bin. While ECE is an intuitive metric for measuring absolute calibration, it fails to capture calibration at the individual prediction level and does not account for the model’s discriminative ability, which refers to how well the model assigns higher confidence to correct predictions over incorrect ones (geng-etal-2024-survey, ). In contrast, AUROC (bradley1997use, ) captures the model’s discriminative ability by computing the probability that a randomly chosen correct prediction is assigned higher confidence than a randomly chosen incorrect one, but it does not measure absolute calibration. Brier Score (brier1950verification, ) quantifies the mean squared difference between predicted confidence and the true binary outcome, capturing both absolute calibration at the individual level and the discriminative ability.
Alternative prompting strategies and setups
To ensure the robustness of our findings, we experiment with a wide range of alternative prompting strategies and setups, and provide the results in Appendix A.2. (1) We explore different confidence expression styles such as providing only a linguistic descriptor without a probability, or outputting a numerical probability directly without binning. (2) Since some reasoning models perform two rounds of inference due to the forced Confidence Reasoning, we also evaluate a similar two-step sequential prompting setup for non-reasoning models. (3) We evaluate more advanced prompting strategies for non-reasoning models, as suggested by prior work, such as Top-K and Multi-step prompting (tian-etal-2023-just, ; xiong2024can, ). (4) While our main experiments use greedy decoding to avoid randomness and ensure reproducibility, we also test sampling-based decoding to confirm that our results generalize across decoding strategies. Across all setups, our findings remain consistent: reasoning models exhibit better calibration than their non-reasoning counterparts.
3.2 Experiment result
We observe that reasoning models consistently outperform their non-reasoning counterparts across all calibration metrics on knowledge-focused datasets (Table 1). Notably, reasoning models exhibit superior calibration even in cases where they underperform non-reasoning models in task accuracy. For example, GLM-Z1-0414 achieves 0.07 lower accuracy than GLM-4-0414 on NonambigQA, yet it obtains a 0.037 lower ECE and a 0.078 higher AUROC. This indicates that accurate confidence estimation is not merely a byproduct of better task performance. Additionally, this highlights a potential motivation for using reasoning models even on simple factual question-answering tasks: although they may not necessarily achieve higher accuracy, they communicate their confidence more reliably.
Interestingly, Qwen3 exhibits a substantial difference in calibration depending on whether it engages in CoT through Thinking or Non-thinking Mode. As Qwen3 is a single checkpoint, this suggests that the calibration observed in reasoning models primarily stems from their slow thinking CoT, rather than some other inherent differences between the models. We further support this claim in Section 4.3, where we show that even non-reasoning models can improve calibration by engaging in slow thinking through in-context learning.
Figure 2: Accuracy (left) and sample frequency (right) across confidence bins for Qwen2.5-32B-Instruct and R1-Distill-Qwen-32B on TriviaQA.
Figure 2 (left) illustrates the superior calibration of R1-Distill-Qwen over Qwen2.5-Instruct on TriviaQA. Surprisingly, for confidence levels above 55%,55555% corresponds to the “Better than even (0.5–0.6)” bin. Throughout the paper, we denote each bin by the average value of its numerical confidence range for brevity. R1-Distill-Qwen exhibits near-perfect calibration, where its estimated confidence closely matches the actual accuracy (i.e., points lie near the line $y=x$, where $y$ is accuracy and $x$ is confidence). Plotting the sample frequency of the confidence bins (right of Figure 2) reveals that Qwen2.5-Instruct is highly overconfident, estimating either the 85% or 95% confidence bin for more than 80% of the total samples. In contrast, R1-Distill-Qwen produces estimations across a relatively diverse range of confidence levels, indicating less tendency to be overconfident. Nonetheless, there is room for improvement for R1-Distill-Qwen as well, as it rarely estimates confidence lower than 55%.
Reasoning models are also better calibrated on reasoning-intensive datasets (Table 2), showing that our findings generalize to more complex tasks. However, we observe a few cases where non-reasoning models outperform their reasoning counterparts in AUROC; specifically, GLM-4-0414 on both Math subsets, and EXAONE-3.5-Instruct on the Non-Math subset of SuperGPQA. Upon closer inspection, we find that the root cause lies in the multiple-choice format of MMLU-Pro and SuperGPQA, which implicitly provides models with cues about their correctness based on the available options. Non-reasoning models, exhibiting overconfidence, tend to select the 95% confidence bin when their answers appear among the multiple-choice options, but when they do not, the models select the closest available option and assign a lower confidence of 85%. Relying on just two confidence bins, reinforced by implicit cues from the multiple-choice options, gives non-reasoning models an unearned advantage in discriminative power, which inflates their AUROC scores. Nonetheless, the non-reasoning models receive worse Brier Scores, as the metric, unlike AUROC, directly penalizes mismatches between predicted confidence and actual outcomes.
Table 2: Benchmark results on reasoning-intensive datasets. The full result with ECE and Accuracy is available in Appendix A.3.
SuperGPQA MMLU-Pro
Math Non-Math Math Non-Math
Model Brier $\downarrow$ AUROC $\uparrow$ Brier $\downarrow$ AUROC $\uparrow$ Brier $\downarrow$ AUROC $\uparrow$ Brier $\downarrow$ AUROC $\uparrow$
Qwen-2.5-32B
Qwen2.5-Inst. 0.295 0.621 0.416 0.522 0.170 0.806 0.283 0.636
[…]
GLM-4-32B-0414
GLM-4-0414 0.256 0.681 0.459 0.516 0.131 0.794 0.282 0.626
[…]
EXAONE-3.5-32B
EXAONE-3.5-Inst. 0.345 0.590 0.441 0.567 0.243 0.676 0.324 0.590
[…]
Qwen3-32B
Qwen3 Non-thinking 0.296 0.588 0.440 0.552 0.155 0.708 0.285 0.632
[…]
4 Analysis: slow thinking enables accurate confidence adjustments
In this section, we present an in-depth analysis showing that the enhanced calibration of reasoning models stems from slow thinking, which enables LLMs to dynamically adjust their confidence throughout the course of reasoning, as illustrated in Figure 1. First, we show that reasoning models gradually produce more accurate confidence estimates as their CoT unfolds, whereas non-reasoning models exhibit no such trend (Section 4.1). We then perform an ablation study, systematically removing components of the CoT, and identify that slow thinking behaviors, specifically exploring alternatives and backtracking, are key contributors to improved calibration (Section 4.2). Finally, we demonstrate that simply prompting non-reasoning models to perform slow thinking through in-context learning yields similar benefits, supporting the notion that these gains are driven by slow thinking itself rather than other inherent differences (Section 4.3).
4.1 (Only) Reasoning models gradually get better calibrated as CoT progresses
Figure 3: Relative change in Brier Score as CoT progresses on NonambigQA. Non-reasoning models are represented by triangles ($\triangle$), and reasoning models by circles ($\bullet$), with each model pair shown in matching colors.
[…]
Figure 3 shows the relative change in Brier Score compared to the initial value at 0 token length (i.e., before any CoT reasoning) on NonambigQA. We observe that all three reasoning models show progressively better calibration with the slope trending downwards. Surprisingly, for non-reasoning models, the Brier Score is the lowest without any CoT, and calibration worsens by the end of the CoT, with no consistent trend observed in between.
To quantify this trend across both datasets and all three metrics, we fit a linear model to each calibration metric over CoT progress and report the slope and its statistical significance across both datasets (Table 3). For reasoning models, we observe strong linear relationships in most cases (except for Qwen3 Thinking’s ECE) with fitted slopes indicating steadily improving calibration over the course of the CoT. Meanwhile, non-reasoning models generally fail to yield statistically significant linear trends, and in the some cases where a statistically significant fit is observed, the slope surprisingly indicates worsening calibration.
Table 3: Slopes of calibration metrics measured over CoT progress. Bold Blue indicates a steady improvement in calibration as CoT progresses, and Bold Red indicates a steady degradation. Bold colors are shown only when the trend is statistically significant ($p<0.05$).
TriviaQA NonambigQA
Model Acc. ECE $\downarrow$ Brier $\downarrow$ AUROC $\uparrow$ Acc. ECE $\downarrow$ Brier $\downarrow$ AUROC $\uparrow$
Qwen2.5-32B
Qwen2.5-Inst. 0.000 0.004 0.000 0.002 0.002 0.002 0.001 0.003
[…]
GLM-4-32B-0414
GLM-4-0414 0.007 -0.003 -0.003 0.000 0.005 0.002 -0.001 0.003
[…]
EXAONE-3.5-32B
EXAONE-3.5-Inst. -0.001 0.001 0.000 0.002 -0.001 0.004 0.003 -0.001
[…]
Qwen3-32B
Qwen3 Non-thinking 0.003 0.004 0.001 -0.001 0.004 0.002 0.002 -0.004
[…]
4.2 Ablation study: exploring alternatives and refining matters the most
To determine which aspects are the most responsible for calibration gains, we analyze R1-Distill-Qwen-32B by removing individual components from its CoT. Specifically, we use the model’s original CoTs on the knowledge-focused datasets, TriviaQA and NonambigQA, and re-prompt the model with targeted components removed. We ablate the following:
  • Confidence Reasoning: We retain only the portion of the CoT that appears before the Confidence Reasoning force prompt (“Okay, now let’s assess my overall thinking process …”) described in Section 3.1, thereby ablating the effect of explicit reasoning about confidence.
  • Epistemic Markers: Reasoning models frequently generate epistemic phrases that signal uncertainty, such as “I think” or “maybe”, which could affect the confidence estimation. Therefore, we remove or paraphrase such phrases, while preserving the rest of the content as closely as possible to the original.
  • Non-linear Reasoning: Reasoning models follow complex non-linear reasoning paths, including behaviors such as exploring alternatives, refining, and backtracking. We prune these non-linear traces, retaining only the parts that directly support the model’s final answer and thus making the reasoning path linear.
We also report the No CoT setting as a lower bound for reference, in which the model’s reasoning is terminated immediately by prompting “</think>”.
For both Epistemic Markers and Non-linear Reasoning removal, we use GPT-4.1 (openai_gpt41, ) guided by detailed instructions and three-shot demonstrations created by the authors. Only the portion of the CoT before the forced Confidence Reasoning prompt is provided. For all three settings, the authors manually inspect 100 examples using predefined criteria, observing that more than 90% meet the intended requirements for each setting. Additional details on the experimental setup, including the full prompts used with GPT-4.1, the manual inspection process, and before-and-after examples of the CoTs, are provided in Appendix B.3.
Table 4: Ablation study on the CoT of R1-Distill-Qwen-32B. The “–” symbol indicates that the corresponding component is removed from the original CoT. The No CoT setting is included as a lower-bound reference.
TriviaQA NonambigQA
Method Acc. ECE $\downarrow$ Brier $\downarrow$ AUROC $\uparrow$ Acc. ECE $\downarrow$ Brier $\downarrow$ AUROC $\uparrow$
[…]
– Confidence Reason. 0.722 0.063 0.166 0.763 0.494 0.187 0.247 0.718
– Epistemic Markers 0.726 0.109 0.154 0.837 0.500 0.262 0.263 0.784
– Non-linear Reason. 0.731 0.161 0.179 0.734 0.500 0.341 0.320 0.728
[…]
Table 4 presents the ablation results. First, we observe that Confidence Reasoning has only a minor effect on calibration, suggesting that the model’s ability to express confidence accurately primarily stems from slow thinking about the question itself, rather than explicitly reasoning about its own confidence. Interestingly, removing Epistemic Phrases leads to a noticeable degradation in ECE, but results in improved AUROC. Upon further investigation, we find that the model becomes overconfident, yet surprisingly retains its ability to discriminate between correct and incorrect answers. Specifically, on TriviaQA, the model tends to predict either 95% or 65% confidence almost exclusively, as opposed to the more varied distribution in the original (see right of Figure 2). Despite this overconfidence, predictions at 95% are generally correct, while those at 65% are not, preserving the model’s discriminative capacity. Restricting predictions to just two confidence bins, instead of distributing across a wider range, results in an advantage over the original model in AUROC.
Finally, our results show that Non-linear Reasoning has the greatest impact on calibration. When the reasoning path is constrained to a linear trajectory, all three metrics degrade, indicating a decline not only in absolute calibration but also in the model’s ability to discriminate between correct and incorrect answers. Overall, our ablation results align with the observations in Section 4.1, indicating that as the CoT unfolds, the model gains more opportunity to express epistemic phrases, consider alternatives, and refine its reasoning which all contributes to improved calibration.
4.3 Non-reasoning models also better express their confidence with slow thinking
So far, we have shown that reasoning models are better calibrated (Section 3), that their calibration improves as the CoT unfolds (Section 4.1), and identified components in slow thinking that contribute to this improvement (Section 4.2). Yet an important question remains: Are calibration gains from slow thinking exclusive to reasoning models that went through extensive fine-tuning or reinforcement learning with specialized reasoning data?
In this section, we demonstrate that non-reasoning models also “better express their confidence” by simply engaging in slow thinking via in-context learning without any parameter updates. Specifically, we collect three random examples of R1-Distill-Qwen-32B slow thinking on held-out TriviaQA samples, and use them as few-shot exemplars to prompt non-reasoning models to reason in a similar manner.
Table 5 shows that non-reasoning models also consistently exhibit improved calibration across both datasets and all three metrics when prompted to engage in slow thinking. This highlights that the calibration gains in reasoning models, stem from the act of slow thinking itself rather than from inherent properties exclusive to reasoning models. This claim is further supported by our findings with Qwen3 in Section 3, where the model demonstrates significantly better calibration when performing CoT in Thinking mode compared to Non-Thinking mode.
Table 5: Benchmark results on non-reasoning models prompted to engage in slow thinking via in-context learning.
TriviaQA NonambigQA
Model Acc. ECE $\downarrow$ Brier $\downarrow$ AUROC $\uparrow$ Acc. ECE $\downarrow$ Brier $\downarrow$ AUROC $\uparrow$
Qwen2.5-32B-Inst. 0.709 0.135 0.192 0.714 0.511 0.309 0.321 0.658
[…]
GLM-4-32B-0414 0.783 0.150 0.180 0.631 0.596 0.338 0.347 0.573
[…]
EXAONE-32B-Inst. 0.728 0.118 0.183 0.712 0.519 0.326 0.331 0.670
[…]
5 Discussion
Does forcing longer CoT lead to improved calibration?
Our findings in Section 4.1 suggest that reasoning models become better calibrated the longer they engage in CoT. This prompted us to test whether calibration can be further improved by forcing reasoning models to slow think longer. Specifically, we apply budget forcing (muennighoff2025s1, ) by appending “Wait,” to the end of the CoT and conducting additional rounds of inference. Figure 4 (left) of Appendix A.1 shows that budget forcing does not necessarily lead to further improvements in calibration, suggesting that calibration gains likely stem from the quality rather than the quantity of slow thinking.
Calibration scales better with model size in reasoning models
Since our experiments are conducted on 32B-scale models, we further examine how the effects of slow thinking vary across different model scales. Figure 4 (right) of Appendix A.1 shows the relative calibration gain of reasoning models over their non-reasoning counterparts across various model scales. We find that the calibration gap widens as model size increases, suggesting that the benefits of slow thinking become more pronounced in larger, more capable LLMs. This trend is encouraging, as it suggests that slow thinking will become an increasingly effective mechanism for calibration as LLMs continue to improve.
Room for improvement even for reasoning models
Generally, we observe that reasoning models still prefer to express high confidence, assigning values below 55% infrequently, as shown in Figure 2 (right). This tendency is further reflected in the calibration metrics: the lower task accuracy on NonambigQA leads to notably higher ECE and Brier Scores compared to TriviaQA (Table 1). While these observations suggest that the challenge of expressing uncertainty remains, our overall results demonstrate that slow thinking meaningfully improves calibration and represents a valuable direction for future model development.
6 Conclusion
In this work, we present a comprehensive study of reasoning models’ ability to express confidence through their own words. Across diverse datasets and model families, we show that reasoning models are consistently better calibrated than their non-reasoning counterparts. Through detailed analysis, we trace these calibration gains to slow thinking behaviors which allow models to dynamically adjust their confidence throughout the chain-of-thought. Overall, our findings reveal that slow thinking offers more than just improved problem-solving: it also enhances the trustworthiness and reliability of LLMs by enabling them to better “know what they know.”
Acknowledgments
This work was supported by the InnoCORE program of the Ministry of Science and ICT (N10250156, 50%) and the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.2022-0-00113, Developing a Sustainable Collaborative Multi-modal Lifelong Learning Framework, 25%; No.RS-2021-II212068, Artificial Intelligence Innovation Hub, 25%).
Appendix A Additional Experiment Results and Details
A.1 Experiment Results in Discussion
In this section, we present the results referenced in the Discussion (Section 5). Specifically, Figure 4 shows the effect of forcing reasoning models to produce longer CoTs (left) and the effect of model scale on calibration (right).
Figure 4: Relative change in Brier Score on NonambigQA under budget forcing (left) and across different model scales (right).
A.2 Expanded results for robustness
In this section, we explore alternative setups and conditions to assess the robustness of our findings.
A.2.1 Alternative Prompting Method for Non-reasoning Models
Table 6: Benchmark results with alternative prompting method for non-reasoning models.
TriviaQA NonambigQA
Model Acc. ECE $\downarrow$ Brier $\downarrow$ AUROC $\uparrow$ Acc. ECE $\downarrow$ Brier $\downarrow$ AUROC $\uparrow$
Qwen2.5-32B
Qwen2.5-Inst. 0.718 0.129 0.176 0.769 0.517 0.297 0.303 0.720
+Two-turn 0.709 0.135 0.192 0.714 0.511 0.309 0.321 0.658
+Top-K 0.725 0.141 0.168 0.780 0.504 0.327 0.320 0.698
+Multi-step 0.734 0.107 0.177 0.735 0.524 0.278 0.305 0.659
[…]
GLM-4-32B-0414
GLM-4-0414 0.814 0.084 0.137 0.675 0.640 0.246 0.269 0.643
+Two-turn 0.783 0.150 0.180 0.631 0.596 0.338 0.347 0.573
+Top-K 0.764 0.139 0.167 0.737 0.554 0.319 0.318 0.695
+Multi-step 0.798 0.105 0.153 0.652 0.667 0.228 0.262 0.589
[…]
EXAONE-3.5-32B
EXAONE-3.5-Inst. 0.715 0.130 0.178 0.749 0.511 0.302 0.302 0.721
+Two-turn 0.728 0.118 0.183 0.712 0.519 0.326 0.331 0.670
+Top-K 0.705 0.160 0.187 0.771 0.487 0.372 0.360 0.693
+Multi-step 0.685 0.115 0.186 0.728 0.501 0.269 0.286 0.698
[…]
Qwen3-32B
Qwen3 Non-thinking 0.711 0.207 0.230 0.650 0.511 0.403 0.403 0.572
+Two-turn 0.713 0.202 0.226 0.674 0.500 0.402 0.397 0.620
+Top-K 0.701 0.219 0.236 0.637 0.502 0.415 0.410 0.590
+Multi-step 0.711 0.175 0.208 0.737 0.501 0.369 0.367 0.634
[…]
We apply alternative, more advanced prompting strategies to non-reasoning models to give them further advantage. Specifically:
  • Two-turn: Since some reasoning models perform two rounds of inference due to the forced Confidence Reasoning (Section 3), we also evaluate a similar two-step sequential prompting setup for non-reasoning models. Specifically, we prompt the non-reasoning models to engage in Solution Reasoning in the first turn, followed by Confidence Reasoning and Confidence Verbalization in the second.
  • Top-K: Following prior work, we prompt the model to generate $K$ candidate answers (with $K=4$, as used in previous studies), each accompanied by a confidence estimate. The final answer is then selected as the one with the highest confidence.
  • Multi-step: As suggested by prior work, we prompt the model to split its reasoning into $K$ steps (with $K=4$), assessing its confidence after each step, and use the final confidence as the output.
Even when non-reasoning models are given additional support through alternative prompting methods, reasoning models still consistently outperform them (Table 6).
A.2.2 Different Confidence Expression Styles
In this section, we investigate the effect of confidence expression styles. Specifically we test, (1) providing only a linguistic descriptor without a probability (“Almost Certain”), and (2) outputting a numerical probability directly without binning (“0.95”).
Tables 7 and 8 present benchmark results when using only linguistic descriptors and when using numerical probabilities without bins, respectively. For using linguistic descriptors, reasoning models outperform non-reasoning models across all metrics and both datasets. In contrast, when models are prompted to output a numerical probability directly, we observe a general degradation in calibration for both reasoning and non-reasoning models. This finding is consistent with previous work suggesting that continuous probability expressions are less desirable for confidence estimation in LLMs [3]. Despite the inadequacy of this setting for reliable confidence estimation, reasoning models still generally outperform non-reasoning models.
Table 7: Benchmark results when providing only the linguistic descriptor.
TriviaQA NonambigQA
Model Acc. ECE $\downarrow$ Brier $\downarrow$ AUROC $\uparrow$ Acc. ECE $\downarrow$ Brier $\downarrow$ AUROC $\uparrow$
Qwen2.5-32B
Qwen2.5-Inst. 0.714 0.099 0.170 0.756 0.516 0.259 0.279 0.705
[…]
GLM-4-32B-0414
GLM-4-0414 0.820 0.077 0.131 0.713 0.639 0.222 0.252 0.672
[…]
EXAONE-3.5-32B
EXAONE-3.5-Inst. 0.722 0.116 0.165 0.760 0.514 0.295 0.293 0.729
[…]
Qwen3-32B
Qwen3 Non-thinking 0.717 0.169 0.213 0.673 0.510 0.370 0.376 0.600
[…]
Table 8: Benchmark results when directly outputting a numerical probability without binning.
TriviaQA NonambigQA
Model Acc. ECE $\downarrow$ Brier $\downarrow$ AUROC $\uparrow$ Acc. ECE $\downarrow$ Brier $\downarrow$ AUROC $\uparrow$
Qwen2.5-32B
Qwen2.5-Instruct 0.720 0.195 0.219 0.735 0.524 0.374 0.387 0.672
[…]
GLM-4-32B-0414
GLM-4-0414 0.808 0.103 0.148 0.782 0.650 0.217 0.285 0.702
[…]
EXAONE-3.5-32B
EXAONE-3.5-Instruct 0.735 0.163 0.194 0.688 0.526 0.357 0.367 0.647
[…]
Qwen3-32B
Qwen3 Non-Thinking 0.707 0.238 0.264 0.739 0.509 0.379 0.448 0.659
[…]
A.2.3 Sampling Instead of Greedy Decoding
Table 9: Benchmark results with sampling for R1-Distill-Qwen-32B. Result of Qwen2.5-32B-Instruct provided for reference. We conduct five runs and report the average and standard deviation of each metric.
TriviaQA NonambigQA
Model Acc. ECE $\downarrow$ Brier $\downarrow$ AUROC $\uparrow$ Acc. ECE $\downarrow$ Brier $\downarrow$ AUROC $\uparrow$
Qwen2.5-32B-Inst. 0.718 0.129 0.176 0.769 0.517 0.297 0.303 0.720
[…]
While our main experiments use greedy decoding to reduce randomness and minimize computational cost, reasoning models are often recommended to be used with sampling. We therefore test whether they retain their superior calibration under these recommended decoding settings.
Table 9 shows that R1-Distill-Qwen-32B continues to outperform Qwen2.5-32B-Instruct in terms of ECE and Brier Score when using its recommended sampling configuration (temperature = 0.6). While its AUROC is slightly lower than that of Qwen2.5-32B-Instruct, further analysis reveals that this drop is due to R1-Distill-Qwen-32B producing a more diverse range of confidence values—unlike Qwen2.5-32B-Instruct, which mostly outputs just two bins. This diversity imposes a disadvantage in AUROC despite stronger absolute calibration performance.
A.2.4 Bootstrapping
Table 10: Benchmark results under stochastic sampling (five bootstrap resamples). We report the average and standard deviation of each metric.
TriviaQA NonambigQA
Model Acc. ECE $\downarrow$ Brier $\downarrow$ AUROC $\uparrow$ Acc. ECE $\downarrow$ Brier $\downarrow$ AUROC $\uparrow$
Qwen2.5-32B-Inst.
- Average 0.743 0.100 0.158 0.774 0.503 0.306 0.308 0.719
- Standard Deviation 0.019 0.017 0.012 0.022 0.010 0.007 0.005 0.010
[…]
To assess variability, we perform bootstrapping by generating five resampled subsets of 1,000 examples for each dataset and report the standard deviation across runs (Table 10). Due to the scale of our experiments, we do not run this analysis across all models and datasets, as doing so would incur unreasonable computational cost.
A.3 Full Results on SuperGPQA and MMLU-Pro
In this section, we provide the full benchmark results on SuperGPQA (Table 11) and MMLU-Pro (Table 12), including both accuracy and ECE.
Table 11: Full benchmark results on SuperGPQA.
Math Non-Math
Model Acc. ECE $\downarrow$ Brier $\downarrow$ AUROC $\uparrow$ Acc. ECE $\downarrow$ Brier $\downarrow$ AUROC $\uparrow$
Qwen2.5-32B
Qwen2.5-Inst. 0.485 0.230 0.295 0.621 0.333 0.431 0.416 0.522
[…]
GLM-4-32B-0414
GLM-4-0414 0.609 0.208 0.256 0.681 0.402 0.460 0.459 0.516
[…]
EXAONE-3.5-32B
EXAONE-3.5-Inst. 0.343 0.336 0.345 0.590 0.238 0.494 0.441 0.567
[…]
Qwen3-32B
Qwen3 Non-thinking 0.604 0.257 0.296 0.588 0.412 0.442 0.440 0.552
[…]
Table 12: Full benchmark results on MMLU-Pro.
Math Non-Math
Model Acc. ECE $\downarrow$ Brier $\downarrow$ AUROC $\uparrow$ Acc. ECE $\downarrow$ Brier $\downarrow$ AUROC $\uparrow$
Qwen2.5-32B
Qwen2.5-Inst. 0.669 0.094 0.170 0.806 0.560 0.230 0.283 0.636
[…]
GLM-4-32B-0414
GLM-4-0414 0.802 0.067 0.131 0.794 0.637 0.250 0.282 0.626
[…]
EXAONE-3.5-32B
EXAONE-3.5-Inst. 0.598 0.156 0.243 0.676 0.498 0.285 0.324 0.590
[…]
Qwen3-32B
Qwen3 Non-thinking 0.787 0.083 0.155 0.708 0.626 0.249 0.285 0.632
[…]
A.4 Full CoT of the Qualitative Analysis
Figure 5 presents the full, untruncated CoT from R1-Distill-Qwen-32B shown in Figure 1.
Question Who’s the Athenian reformer that set up the council of 500?
Model CoT Okay, so I need to figure out which Athenian reformer set up the council of 500. Hmm, I remember that Athens had several important reformers, like Solon, Cleisthenes, and Pericles. Let me think about each of them. Solon was an early reformer, right? He introduced some democratic elements, like the council of 400, I think. But wait, the question is about the council of 500. So maybe it wasn’t Solon. Then there’s Cleisthenes. I recall that he was a significant figure in the development of Athenian democracy. He reorganized the political system after the overthrow of the tyrants. He introduced the concept of demes, which were local administrative units, and he increased the power of the Assembly. (1) Oh, and he also created the council of 500, which was called the Boule. Each deme would send a certain number of representatives to this council, which would meet regularly and handle day-to-day affairs. That makes sense because the council of 500 was a key part of the democratic reforms in Athens. (2) Pericles came later, and he was more known for his leadership during the Golden Age of Athens, including the construction of the Parthenon and the expansion of democracy, (3) but I don’t think he was the one who established the council of 500. That was already in place before his time. (4) So putting it all together, Cleisthenes is the reformer who set up the council of 500. I’m pretty confident about this because it’s a well-known part of his reforms. I don’t think I’m mixing up any details here. The council of 500 was a major institution in the Athenian democracy, and Cleisthenes is credited with its creation as part of his efforts to decentralize power and make the government more representative.
Figure 5: Full CoT version of Figure 1.
Appendix B Experimental Setup Details
B.1 Experiments Compute Resources
We conduct our experiments on machines equipped with either Nvidia A6000 48GBs or A100 80GBs GPUs. For evaluating 32B-scale models, we use two GPUs. Leveraging vLLM [18] for efficient inference, evaluation of a 32B reasoning model on a single dataset takes approximately one hour on two A6000s.
B.2 Additional Details on Inference Procedure
The full prompt given to the models to perform the three steps—Solution Reasoning, Confidence Reasoning, and Confidence Verbalization—along with the required answer and confidence formatting (“Answer:ANSWER Confidence:CONFIDENCE”) is provided in Listing B.2.
Listing 1: Prompt used throughout our experiments First, reason through the question step by step to arrive at an answer. Then, thoroughly assess your confidence in that answer by evaluating your thinking process so far. Finally, classify your confidence into one of the following classes based on how likely your answer is to be correct: - "Almost no chance" (0.0–0.1) - "Highly unlikely" (0.1–0.2) - "Chances are slight" (0.2–0.3) - "Unlikely" (0.3–0.4) - "Less than even" (0.4–0.5) - "Better than even" (0.5–0.6) - "Likely" (0.6–0.7) - "Very good chance" (0.7–0.8) - "Highly likely" (0.8–0.9) - "Almost certain" (0.9–1.0) Each category reflects the probability that your answer is correct. At the very end of your output, format your answer and confidence as **Answer**: \$ANSWER **Confidence**: \$CLASS where CLASS is one of the names (only the names without the probability ranges) of the classes above.
During our preliminary experiments, we occasionally observed cases—though rare—where models failed to produce the answer and confidence in the required format. To ensure robust extraction, we conduct an additional round of inference: for non-reasoning models, we request only the answer and confidence to be re-output in the specified format; for reasoning models, we first generate up to the </think> token, then replace it with </think>Answer: and perform a second round of inference.
For both reasoning and non-reasoning models, we use greedy decoding, with the maximum token length set to 4096 for knowledge-focused datasets and 8192 for reasoning-intensive datasets. The effect of using sampling is examined in Appendix A.2.
B.3 Additional Details on Ablation Study
In this section, we detail the process for assessing the quality of the CoTs used in the ablation study in Section 4.2, along with before-and-after examples (Listing B.3), and full prompts given to GPT-4.1 (Listing B.3 and Listing B.3). For quality check, we do the following:
  • Confidence Reasoning: The authors manually inspect 100 CoT examples sampled from both datasets we find that only four examples contain explicit reasoning about the model’s own confidence.
  • Epistemic Phrases: The authors inspect 100 samples collected from both datasets and evaluate them based on two criteria: (1) whether epistemic phrases are absent, and (2) whether the remaining content stays true to the original. We find that 91 out of 100 examples satisfy both criteria.
  • Non-linear Reasoning: We again provide detailed instructions along with three-shot examples. The authors then manually inspect 100 samples and evaluate them based on two criteria: (1) whether only the content supporting the final answer remains, and (2) whether the remaining content stays faithful to the original. We find that all 100 examples meet both criteria.
Listing 2: Before-and-after examples of CoT [Original (No Confidence Reasoning)]
Okay, so I need to figure out which country the Order of the Elephant is the highest order of. I’m not too familiar with this, but I’$l$ try to reason it out.
First, I remember that orders are usually honors or decorations given by countries to recognize people for their contributions. The Order of the Elephant sounds like a very prestigious award. I think it’s associated with Denmark because I’ve heard of the Order of the Elephant in the context of Danish honors. But I’m not entirely sure. Wait, maybe it’s from another country. Let me think. I know the Order of the British Empire is from the UK, and the Order of Canada is from Canada. The Order of the Elephant doesn’t ring a bell as being from those places. I think Denmark has a few orders, and the Order of the Elephant might be one of them. I’m pretty sure it’s a Danish order, but I’m not 100% certain. I should also consider if there are other countries with similar-sounding orders. Maybe Sweden or Norway? I don’t recall any orders named Elephant from those countries. Denmark seems more likely because I’ve heard it mentioned in the context of their royal family or state honors. So, putting it all together, I think the Order of the Elephant is the highest order of Denmark. I’m fairly confident, but I’m not 100% sure because I don’t have direct knowledge. I might be mixing it up with another order, but I can’t think of any others right now. [No Epistemic Phrases]
Okay, so I need to figure out which country the Order of the Elephant is the highest order of. I’$l$ try to reason it out.
First, orders are usually honors or decorations given by countries to recognize people for their contributions. The Order of the Elephant sounds like a prestigious award. It is associated with Denmark, and the Order of the Elephant appears in the context of Danish honors. Wait, maybe it’s from another country. The Order of the British Empire is from the UK, and the Order of Canada is from Canada. The Order of the Elephant does not appear to be from those places. Denmark has a few orders, and the Order of the Elephant is one of them. It is a Danish order. Consideration can be given to other countries with similar-sounding orders, such as Sweden or Norway. However, there are no orders named Elephant from those countries. Denmark is more likely, as it is mentioned in the context of their royal family or state honors. So, the Order of the Elephant is the highest order of Denmark. [No Non-linear Reasoning]
Okay, so I need to figure out which country the Order of the Elephant is the highest order of.
Orders are honors given by countries, and the Order of the Elephant is associated with Denmark. It’s mentioned in the context of the Danish royal family or state honors. So, the Order of the Elephant is the highest order of Denmark.
Listing 3: Prompt given to GPT-4.1 to remove Epistemic Phrases I will give you a model’s thinking process to a question. I want you to remove explicit expressions of uncertainty or confidence (like "I’m not sure", "I think", "maybe", "I’m a bit confused", "I’m pretty confident", etc.) in the model’s thinking process. Here are the details: - Preserve the original thinking process and structure — do not delete any logical steps even if they seem mistaken or redundant. - Rephrase uncertain sentences into neutral equivalents whenever possible. For example, change "I’m not sure if he wrote all three books or if someone else took over" into "He either wrote all three books or someone else took over." - Preserve "Wait," whenever it appears. It signals a natural shift or reevaluation, not uncertainty. - Don’t shortcut the reasoning or fix factual errors. - Keep the output as close as possible to the original wording, only editing where necessary to remove uncertainty. - If the original thinking degenerates into repetition, or fails to reach a conclusion, stop the rewrite at the point where the reasoning becomes incomplete or broken. Do not invent, infer, or complete missing reasoning that was not present in the original. - Strictly repeat the first sentence of the model’s thinking process in the output. - Use double linebreaks between paragraphs. Here are some examples:
Listing 4: Prompt given to GPT-4.1 to remove Non-linear Reasoning I will give you a model’s thinking process to a question. I want you to rewrite it into a concise and linear version, meaning the reasoning should move step-by-step directly toward the final answer without backtracking, or unnecessary side exploration. Also remove any expressions of uncertainty or confidence (like "I’m not sure", "I think", "maybe", "I’m a bit confused", "I’m pretty confident", etc.). Here are the details: - Linearize the reasoning — remove backtracking, "Wait," moments, or diversions. - Keep only the details or examples directly supporting the final answer to make it concise. - Remove any explicit expressions of uncertainty or confidence (such as "maybe", "I’m not sure", "I think", "I’m pretty confident", etc.). - Do not fix any factual errors. - Keep the tone casual and natural, like someone logically talking to themselves. - If the original thinking degenerates into repetition, or fails to reach a conclusion, stop the rewrite at the point where the reasoning becomes incomplete or broken. Do not invent, infer, or complete missing reasoning that was not present in the original. - Strictly repeat the first sentence of the model’s thinking process in the output. - Use double linebreaks between paragraphs. Here are some examples:
Appendix C Licenses for existing assets
C.1 Models
  • R1-Distill [4]: MIT
  • QwQ [35]: Apache-2.0
  • OR1-Preview [11]: Apache-2.0
  • Qwen-2.5 [33]: Apache-2.0
  • Qwen-3 [34]: Apache-2.0
  • GLM-4-0414 [39]: MIT
  • GLM-Z1-0414 [39]: MIT
  • EXAONE-3.5 [19]: EXAONE
  • EXAONE-Deep [19]: EXAONE
C.2 Datasets
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We claim that reasoning models are better calibrated, and slow thinking enables this capability. This claim is supported throughout Section 3 and Section 4.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: While we have demonstrated that reasoning models better express their confidence, we acknowledge and discuss that there’s still room for improvement in Section 5.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: Our work does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide detailed explanation regarding our experiment setting in Section 3.1 and Appendix B. Also our code is available for access.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Our code is available at the anonymous link mentioned at the first page of the paper, as well as in the supplementary material.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Our full experimental setting and details are explained throughout Section 3.1 and Appendix B.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: To support our main claim that reasoning models are better calibrated, we perform bootstrapping by generating five resampled subsets and report the standard deviation across runs at Appendix A.2.4.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide details on our compute resources in Appendix B.1.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The authors have reviewed the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Our work does not have negative societal impacts. Throughout the paper, we mention that our findings can lead to development of trustworthy and reliable LLM systems.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We cite all the assets we use in Section 3.1 and list the licenses in Section C.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We do not release new assets other than the experiment code. The documentation for the code is provided in the README.md file which can be accessed via the anonymized URL or from the supplementary material.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Our paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Our paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: We leverage LLMs to evaluate model responses and to create the ablation data, as mentioned in Section 3.1 and Section 4.2.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.